Find the value of m for which lim x->infinity


(x+1)(2x-3)(2-3x)/(4-x+mx³) exists.

1
Expert's answer
2021-11-30T13:09:28-0500

"\lim_{x\to\infty} \frac{(x+1)(2x-3)(2-3x)}{(4-x+mx^3)}\\
=\lim_{x\to\infty} \frac{-6x^3+7x^2+7x-6}{(4-x+mx^3)}\\"

Divide numerator and denominator by "x^3"

"=\lim_{x\to\infty} \frac{-6+\frac{7}{x}+\frac{7}{x^2}-\frac{6}{x^3}}{(\frac{4}{x^3}-\frac{1}{x^2}+m)}\\
=\frac{-6}{m}"

For limit to be exist, "m\neq0"